Find the dimensions of the rectangle that has a length of 3 meters more than its width and a perimeter equal in value to its area?
Found 2 solutions by josgarithmetic, lwsshak3:
Answer by josgarithmetic(39617)

L=3+w and 2w+2L=wL.

Substitute for L in the perimeter&area equation.
2w+2(w+3)=w(w+3)
Go from that.

Answer by lwsshak3(11628)

Find the dimensions of the rectangle that has a length of 3 meters more than its width and a perimeter equal in value to its area?
***
let x=width
x+3=length
..
2(width+length)=2(x+x+3=4x+6=perimeter
width*length=x(x+3)=x^2+3x=area
4x+6=x^2+3x
x^2-x-6=0
(x-3)(x+2)=0
x=3
x+3=6
dimensions of the rectangle:
width=3
length=6